25. A partly full paint can has 2.75 liters of paint left in it. What is volume of the paint in cubic centimeters? (3)

A) 2.75 x 10-3 cm3 C) 275 x 10-3 cm3 E) 27.5 x 103 cm3

B) 2.75 cm3 D) 2.75 x 103 cm3 



1
Expert's answer
2021-06-13T07:09:46-0400

25)


"2.75\ L\cdot\dfrac{0.001\ m^3}{1\ L}\cdot\dfrac{10^6\ cm^3}{1\ m^3}=2750\ cm^3=2.75\cdot10^3\ cm^3."

Therefore, the correct answer is (D) 2.75 x 103 cm3